# Cutoff for the Swendsen–Wang dynamics on the complete graph

### v1

@@ 4 Convergence to the neighborhood of the majority vector
By Fact 2.5, $|F^{\prime}(a)|<1$, so for small enough $\varepsilon$, $\max_{|x-a|<\varepsilon}|F^{\prime}(x)|<1$. Also, since $\alpha_{1}(X_{0})\in(a-\varepsilon,a+\varepsilon)$ by assumption, Facts 2.5 and 2.6 imply that so does $F(\alpha_{1}(X_{t-2}))$. Hence, the Taylor expansion of the function $F$ about $F(\alpha_{1}(X_{t-2}))$ yields that for a suitable $\rho\in(0,1)$
@@
Iterating this argument, we deduce that
@@
For this, let $K>0$ and let us define the sequence
$r_{i}=K\big{(}1-{\rho}/{2}\big{)}^{t-i}\big{(}1+\rho\big{)}^{t-i}$ (14)
By Lemma 4.3, we can choose $K$ large enough such that
@@
If $\xi_{i}<r_{i}\sqrt{n}$ for all $i\leq t$, then
for large enough $C$. It then follows from a union bound that
@@
For ease of notation, set $|F^{\prime}(a)|=\rho<1$. By Fact 2.7, there exists $\Delta\geq|F^{\prime\prime}(x)|$ for all $x\in(a-\varepsilon,a+\varepsilon)$. Let $d_{t}=|a-F^{t}(x_{0})|$. The Taylor expansion of $F$ about $a$ implies that
$\rho d_{t-1}-\frac{\Delta}{2}d_{t-1}^{2}\leq d_{t}\leq\rho d_{t-1}+\frac{\Delta}{2}d_{t-1}^{2}.$ (15)
@@
$\rho^{T}d_{0}-\frac{\Delta}{2\rho}\sum_{i=1}^{T}\rho^{i}d_{T-i}^{2}\leq d_{T}\leq\rho^{T}d_{0}+\frac{\Delta}{2\rho}\sum_{i=1}^{T}\rho^{i}d_{T-i}^{2}.$
Since $d_{t}\leq\varepsilon$ for $t\geq 0$, for $\varepsilon$ small enough, we can deduce from (15), that there exists a constant $\varepsilon_{0}$ such that $\frac{\Delta\varepsilon}{2}\leq\varepsilon_{0}<1-\rho$ such that $d_{t}\leq(\rho+\varepsilon_{0})d_{t-1}$, and thus
$d_{T}\leq\frac{|x_{0}-a|\rho^{\gamma}}{\sqrt{n}}+\frac{\Delta(\rho+\varepsilon_{0})^{2T}}{2\rho}\sum_{i=1}^{T}\Big{(}\frac{\rho}{(\rho+\varepsilon_{0})^{2}}\Big{)}^{i}\leq\frac{\varepsilon\rho^{\gamma}}{\sqrt{n}}+\frac{\Delta(\rho+\varepsilon_{0})^{2T}}{2\rho}\sum_{i=1}^{T}\Big{(}\frac{\rho}{(\rho+\varepsilon_{0})^{2}}\Big{)}^{i}.$
We can choose $\varepsilon_{0}$ so that $\rho<(\rho+\varepsilon_{0})^{2}$. Then, $(\rho+\varepsilon_{0})^{2T}=o(1/\sqrt{n})$ and $\sum_{i=1}^{T}\Big{(}\frac{\rho}{(\rho+\varepsilon_{0})^{2}}\Big{)}^{i}=O(1)$, and since $\rho^{T}=\frac{\rho^{\gamma}}{\sqrt{n}}$, we obtain
$\frac{d_{0}\rho^{\gamma}}{\sqrt{n}}-o\big{(}\frac{1}{\sqrt{n}}\big{)}\leq d_{T}\leq\frac{d_{0}\rho^{\gamma}}{\sqrt{n}}+o\big{(}\frac{1}{\sqrt{n}}\big{)}.$
@@ 4.1 Mixing time lower bound
If $X\sim\mu$, then $\|\alpha(X)-m\|_{\infty}\leq A/{\sqrt{n}}$ with probability at least 3/43/43 / 4 for a suitable constant $A>0$.
Very similar results are available in the literature, but we were not able to find this particular variant of the concentration bound, so, for completeness, we provide a proof of Lemma 4.4 in Appendix B.
@@
By Lemma 4.4, if $X\sim\mu$, then $\|\alpha(X)-m\|_{\infty}\leq A/{\sqrt{n}}$ with probability at least 3/43/43 / 4 for a suitable constant $A$. Now, take $X_{0}$ to be a configuration where all vertices are assigned the same spin. By Lemmas 4.1 and 4.2, there exists $\gamma$ such that after $T=c(\beta,q)\log n+\gamma$ steps, with probability 3/4$\geq$ 3/4≥ 3 / 4, we have $\|\alpha(X_{T})-m\|_{\infty}>\frac{A}{\sqrt{n}}$. This implies that $\|X_{T}-\mu\|_{\textsc{tv}}>1/4$ and thus that the SW dynamics has not mixed after $T$ steps. ∎
@@ 5 Spin fraction coalescence in partition
Let $\{V_{1},\dots,V_{q}\}$ be a $\lambda$-partition of $V$ and suppose $X_{t}$ is such that $\|\alpha(X_{t})-m\|_{\infty}\leq\varepsilon$ for a sufficiently small $\varepsilon>0$. Then, there exists $\gamma_{1}(\beta,q)$ such that for any $r\geq\gamma_{1}(\beta,q)$ and any $k\in[q]$
@@
Since $\beta>q$ and $a>1/q$ by Fact 2.5, $a\beta>1$ and $b\beta<1$. By assumption $\|\alpha(X_{t})-m\|_{\infty}\leq\varepsilon$, so when $\varepsilon$ is small enough, exactly one color class in the percolation step of the SW dynamics is supercritical, and the remaining $q-1$ are subcritical.
@@
$\Pr\big{[}R_{1}\geq\gamma_{0}n+\gamma_{2}\sqrt{\gamma_{1}n}\big{]}\leq\Pr\big{[}R\geq\gamma_{0}n+{\gamma_{2}}\sqrt{\gamma_{1}n}\big{]}\leq\Pr\Big{(}R\geq\mathrm{E}[R]+\frac{\gamma_{2}}{2}\sqrt{\gamma_{1}n}\Big{)}\leq\frac{4}{\gamma_{2}^{2}}.$
@@
$Q_{j}^{(i)}=\begin{cases}|C_{j}\cap V_{1}|&\text{if}~C_{j}\text{ is colored }i;~\text{and}\\ 0&\text{otherwise}.\\ \end{cases}$
@@
$\Pr\left(\big{|}|C_{1}\cap V_{1}|-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1,1}(X_{t})|V_{1}|\big{|}\geq\gamma_{4}\sqrt{n}\right)\leq c_{1}\exp{(-c_{2}\gamma_{4}^{2})}$
If $||C_{1}\cap V_{1}|-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1,1}(X_{t})|V_{1}|\big{|}<\gamma_{4}\sqrt{n}$, and $|Z_{j}-E[Z_{j}]|<\gamma_{3}\sqrt{n}$ for all $j\in[q]$, we have
@@
$\alpha_{1,1}(X_{t+1})|V_{1}|\geq G(\alpha_{1}(X_{t}),\alpha_{1,1}(X_{t+1}))n-(\gamma_{3}+\gamma_{4})\sqrt{n},$
@@
$\leq c_{1}\exp(-c_{2}\gamma_{4}^{2})+2\exp\Big{(}\frac{-\gamma_{3}^{2}}{\gamma_{0}+\gamma_{2}\sqrt{\gamma_{1}/n}}\Big{)}+\frac{4}{\gamma_{2}^{2}}$
@@
Recall that $d_{t+1}^{(1)}=|\alpha_{1}(X_{t+1})-\alpha_{1,1}(X_{t+1})|$ and
$\hat{\xi}_{t+1}=\Big{|}d_{t+1}^{(1)}-\Big{(}1-\frac{1}{q}\Big{)}\theta(\beta\alpha_{1}(X_{t}))d_{t}\Big{|}.$
@@
$\Pr\Big{(}\hat{\xi}_{t+1}\geq\frac{r}{\sqrt{n}}\Big{)}\leq e^{-\Omega(r)}+O(r^{-2}),$
@@
Let $\hat{K}>0$ and define the sequence $\hat{r}_{i}=\hat{K}\big{(}1-{\varepsilon}/{2}\big{)}^{t-i}\big{(}1+\varepsilon\big{)}^{t-i}.$ By Lemma 5.1, for $\hat{K}$ large enough
@@
$\sum_{i=1}^{t}(1-\varepsilon)^{t-i}\hat{\xi}_{i}^{(k)}\leq\frac{\hat{K}}{\sqrt{n}}\sum_{i=1}^{t}(1-\varepsilon)^{t-i}\big{(}1-{\varepsilon}/{2}\big{)}^{t-i}\big{(}1+\varepsilon\big{)}^{t-i}\leq\frac{C}{2\sqrt{n}},$
@@
$\Pr\Big{(}\sum_{i=1}^{t}\hat{\xi}_{i}^{(k)}(1-\varepsilon)^{t-i}\geq\frac{C}{2\sqrt{n}}\Big{)}\leq\sum_{i=1}^{t}\Pr\Big{(}\hat{\xi}_{i}^{(k)}>\frac{\hat{r}_{i}}{\sqrt{n}}\Big{)}\leq\frac{\delta}{q}.$
[…]
@@ 6 Projection chain
[…]
$\Pr(|X-Y|=\sqrt{n}\cdot(c_{1},c_{2},\dots,c_{q}))\geq 1-\gamma.$
@@
We couple $A_{t+1}$ and $A_{t+1}^{\prime}$ as follows. The percolation step is done independently in both copies. In the coloring step, first, the same color is assigned to the largest connected component in both chains, and all other connected components with two or more vertices from both chains are assigned spins uniformly at random independently. In addition, in each $V_{i}$, if one copy has more isolated vertices than the other, the excess of isolated vertices in that copy are assigned spins independently. The coloring of the remaining isolated vertices will be coupled to fix the potential discrepancies created in each $V_{i}$ by these steps as we described next.
For $i,j\in[q]$, let $\hat{a}_{ij}$, $\hat{a}_{ij}^{\prime}$ be the number of vertices assigned color $j$ in $V_{i}$ so far (i.e., excluding the isolated vertices). We claim that there exists a constant $C>0$ such that with probability $\Omega(1)$ it holds for all $i,j\in[q]$ that
@@
To prove this, let $\{C_{k}\}_{k\geq 1}$ be the connected components in the first chain after the percolation step; assume $C_{1}$ is the largest component. Let $C(v)$ be the connected component of vertex $v$ after the percolation step. Note that
$\sum_{k\geq 2}|C_{k}\cap V_{i}|^{2}=\sum_{k\geq 2}\sum_{v\in C_{k}\cap V_{i}}\sum_{w\in C_{k}\cap V_{i}}1=\sum_{k\geq 2}\sum_{v\in C_{k}\cap V_{i}}|C_{k}\cap V_{i}|=\sum_{k\geq 2}\sum_{v\in C_{k}\cap V_{i}}|C(v)\cap V_{i}|\leq\sum_{v\in V_{i}\setminus C_{1}}|C(v)|.$
@@
A union bound the implies that $\sum_{k\geq 2}|C_{k}\cap V_{i}|^{2}=O(|V_{i}|)$ for all $i\in[q]$ with probability at least 99/10099/10099 / 100.
Let $\hat{n}_{i}$ be the number of vertices in $V_{i}$ excluding those in $C_{1}$ and the isolated vertices. If $C_{1}$ is assigned spin $s$, then conditioning on $\sum_{k\geq 2}|C_{k}\cap V_{i}|^{2}<100qA_{1}|V_{i}|$, by Hoeffding’s inequality, we deduce that for a large enough constant $A_{2}>0$, with probability at least $\frac{99}{100q^{2}}$, for any $i,j\in[q]$ with $j\neq s$, it holds that
@@
Similarly, we can get an analogous bound for $\hat{a}_{ij}^{\prime}$, and by a union bound (18) holds for all $i,j\in[q]$ such that $j\neq s$ with probability at least 98/10098/10098 / 100.
By Lemma 2.3 and assumption A2, for a suitable constant $A_{2}>0$, we have for any $i\in[q]$ that
@@
Therefore, by a union bound, for all $i\in[q]$
$\Big{|}\hat{a}_{ik}-|C_{1}\cap V_{i}|-\frac{\hat{n}_{k}}{q}\Big{|}$ $\leq(A_{2}+A_{3})\sqrt{n},~\text{and}$
$\Big{|}\hat{a}_{ik}^{\prime}-|C_{1}^{\prime}\cap V_{i}|-\frac{\hat{n}_{k}}{q}\Big{|}$ $\leq(A_{2}+A_{3})\sqrt{n}$
with probability at least 99/10099/10099 / 100, which implies that (18) holds for all $i\in[q]$ for $j=k$ with that probability.
@@
To formalize this, let $P_{t}(r)$ be the event at time $t$, we have $\|\alpha(X_{t})-m\|_{\infty}\leq r/\sqrt{n}$, $\max_{k\in[q]}|\alpha_{1}(X_{t})-\alpha_{1,k}|\leq r/\sqrt{n}$, and similarly for $Y_{t}$.
@@
For $\delta\in(0,1)$, there exists $r=r(\delta)>0$ such that $\Pr(P_{t+1}(r)\mid P_{t}(r))\geq\delta$.
@@
By assumption, $\|\alpha(X_{t})-m\|_{\infty}\leq r/\sqrt{n}$. Let us assume that $\alpha_{1}(X_{t})\geq a$ (the case when $\alpha_{1}(X_{t})<a$ is analogous). By Fact 2.6, $F(\alpha_{1}(X_{t}))\leq F(a+r/\sqrt{n})$. Moreover, by Fact 2.5 and the mean value theorem, there exists a constant $\eta\in(0,1)$ such that $F(a+r/\sqrt{n})-a=\eta r/\sqrt{n}$ which implies that $(a+r)-F(a+r)=(1-\eta)r/\sqrt{n}$. Hence, in order for $\|\alpha(X_{t+1})-m\|_{\infty}\geq r/\sqrt{n}$, we would need $\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq\frac{(1-\eta)r}{\sqrt{n}}$. From Lemma 4.3, for any large enough constant $r$, we have
$\Pr\Big{(}\|\alpha(X_{t+1})-m\|_{\infty}\geq\frac{r}{\sqrt{n}}\Big{)}\leq\Pr\Big{(}\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq\frac{(1-\eta)r}{\sqrt{n}}\Big{)}\leq O(r^{-2});$

### v2

@@ 4 Convergence to the neighborhood of the majority vector
Since from (10) we know that $F(\alpha_{1}(X_{t-2}))\in(a-2\varepsilon,a+2\varepsilon)$. Hence, the Taylor expansion of the function $F$ about $F(\alpha_{1}(X_{t-2}))$ yields that
@@
Iterating this argument, making use of  (10) in each step, we deduce that
@@
When $\xi_{i}<r_{i}\sqrt{n}$ for all $i\leq t$, then
for large enough $C$. By Lemma 4.3, we can choose $K$ large enough such that
@@
It then follows from a union bound that
@@
For ease of notation, set $|F^{\prime}(a)|=\eta<1$. By Fact 2.7, there exists $\Delta\geq|F^{\prime\prime}(x)|$ for all $x\in(a-\varepsilon,a+\varepsilon)$. Let $d_{t}=|a-F^{t}(x_{0})|$. The Taylor expansion of $F$ about $a$ implies that
$\eta d_{t-1}-\frac{\Delta}{2}d_{t-1}^{2}\leq d_{t}\leq\eta d_{t-1}+\frac{\Delta}{2}d_{t-1}^{2}.$ (18)
@@
$\eta^{T}d_{0}-\frac{\Delta}{2\eta}\sum_{i=1}^{T}\eta^{i}d_{T-i}^{2}\leq d_{T}\leq\eta^{T}d_{0}+\frac{\Delta}{2\eta}\sum_{i=1}^{T}\eta^{i}d_{T-i}^{2}.$
Since $d_{t}\leq\varepsilon$ for $t\geq 0$, for $\varepsilon$ small enough, we can deduce from (18) that
$\big(\eta-\frac{\Delta\varepsilon}{2}\big)d_{t-1}\leq d_{t}\leq\big(\eta+\frac{\Delta\varepsilon}{2}\big)d_{t-1},$
and thus
$\frac{d_{0}\eta^{\gamma}}{\sqrt{n}}-\frac{\Delta d_{0}^{2}(\eta-\frac{\Delta\varepsilon}{2})^{2T}}{2\eta}\sum_{i=1}^{T}\Big(\frac{\eta}{(\eta-\frac{\Delta\varepsilon}{2})^{2}}\Big)^{i}\leq d_{T}\leq\frac{d_{0}\eta^{\gamma}}{\sqrt{n}}+\frac{\Delta d_{0}^{2}(\eta+\frac{\Delta\varepsilon}{2})^{2T}}{2\eta}\sum_{i=1}^{T}\Big(\frac{\eta}{(\eta+\frac{\Delta\varepsilon}{2})^{2}}\Big)^{i}.$
For $\varepsilon$ small enough, $\eta>(\eta+\frac{\Delta\varepsilon}{2})^{2}>(\eta-\frac{\Delta\varepsilon}{2})^{2}$, and thus there exist constants $\gamma_{1}=\gamma_{1}(\eta,\Delta,\varepsilon)>0$ and $\gamma_{2}=\gamma_{2}(\eta,\Delta,\varepsilon)>0$ such that
$d_{T}\leq\frac{d_{0}\eta^{\gamma}}{\sqrt{n}}+\frac{\Delta d_{0}^{2}\gamma_{1}}{2\eta}\frac{\eta^{\gamma}}{\sqrt{n}}=\frac{d_{0}\eta^{\gamma}(1+\frac{\Delta\varepsilon\gamma_{1}}{2\eta})}{\sqrt{n}},$
and
$d_{T}\geq\frac{d_{0}\eta^{\gamma}}{\sqrt{n}}-\frac{\Delta d_{0}^{2}\gamma_{2}}{2\eta}\frac{\eta^{\gamma}}{\sqrt{n}}=\frac{d_{0}\eta^{\gamma}(1-\frac{\Delta\varepsilon\gamma_{2}}{2\eta})}{\sqrt{n}},$
as claimed. ∎
@@ 4.2 Mixing time lower bound
If $X\sim\mu$, then $\|\alpha(X)-m\|_{\infty}\leq A/{\sqrt{n}}$ with probability at least $3/4$ for a suitable constant $A>0$.
Similar results are available in the literature, but we were not able to find this particular variant, so, for completeness, we provide a proof of Lemma 4.4 in Appendix B.
@@
By Lemma 4.4, if $X\sim\mu$, then $\|\alpha(X)-m\|_{\infty}\leq A/{\sqrt{n}}$ with probability at least $3/4$ for a suitable constant $A$. Now, take $X_{0}$ to be a configuration where all vertices are assigned the same spin. By Lemmas 4.1 and 4.2, there exists $\gamma$ such that after $T=c(\beta,q)\log n+\gamma$ steps, with probability $\geq 3/4$, we have $\|\alpha(X_{T})-m\|_{\infty}>\frac{A}{\sqrt{n}}$. This implies that $\|X_{T}-\mu\|_{\textsc{tv}}>1/4$ and thus that the SW dynamics has not mixed after $T$ steps. ∎
@@ 5 Spin fraction coalescence in partition
Let $\{V_{1},\dots,V_{q}\}$ be a $\lambda$-partition of $V$ and suppose $X_{t}$ is such that $\|\alpha(X_{t})-m\|_{\infty}\leq\varepsilon$ for a sufficiently small $\varepsilon>0$. Then, there exists $\gamma_{1}(\beta,q)>0$ such that for any $r\geq\gamma_{1}(\beta,q)$ and any $k\in[q]$
@@
Since $\beta>q$ and $a>1/q$ by Fact 2.5, $a\beta>1$ and $\frac{\beta(1-a)}{q-1}<1$. By assumption $\|\alpha(X_{t})-m\|_{\infty}\leq\varepsilon$, so when $\varepsilon$ is small enough, exactly one color class in the percolation step of the SW dynamics is supercritical, and the remaining q1q-1 are subcritical.
@@
$\Pr\big[R_{1}\geq\gamma_{0}n+\gamma_{2}\sqrt{\gamma_{1}n}\big]\leq\Pr\big[R\geq\gamma_{0}n+{\gamma_{2}}\sqrt{\gamma_{1}n}\big]\leq\Pr\Big(R\geq\mathrm{E}[R]+\frac{\gamma_{2}}{2}\sqrt{\gamma_{1}n}\Big)\leq\frac{4}{\gamma_{2}^{2}}.$ (19)
@@
[…]
@@
[…]
If $C_{1}\cap V_{1}|-\theta(\beta\alpha_{1}(X_{t}))\alpha_{1,1}(X_{t})|V_{1}|\big|<\gamma_{4}\sqrt{n}$, and $|Z_{j}-E[Z_{j}]|<\gamma_{3}\sqrt{n}$ for all $j\in[q]$, we have
@@
$\alpha_{1,1}(X_{t+1})|V_{1}|\geq G(\alpha_{1}(X_{t}),\alpha_{1,1}(X_{t+1}))|V_{1}|-(\gamma_{3}+\gamma_{4})\sqrt{n},$
@@
$\leq c_{1}\exp(-c_{2}\gamma_{4}^{2})+2q\exp\Big(\frac{-\gamma_{3}^{2}}{\gamma_{0}+\gamma_{2}\sqrt{\gamma_{1}/n}}\Big)+\frac{4q}{\gamma_{2}^{2}}$
@@
$d_{t+1}^{(1)}$ $=|\alpha_{1}(X_{t+1})-\alpha_{1,1}(X_{t+1})|,$
$\hat{\xi}_{t+1}^{(1)}$ $=\Big|d_{t+1}^{(1)}-\Big(1-\frac{1}{q}\Big)\theta(\beta\alpha_{1}(X_{t}))d_{t}\Big|,~\text{and}$
@@
$\Pr\Big(\hat{\xi}_{t+1}^{(1)}\geq\frac{r}{\sqrt{n}}\Big)\leq e^{-\Omega(r)}+O(r^{-2}),$
@@
Let $\hat{K}>0$, $\hat{\rho}=1/q$, and define the sequence $\hat{r}_{i}=\hat{K}\big(1-{\hat{\rho}}/{2}\big)^{t-i}\big(1+\hat{\rho}\big)^{t-i}.$ By Lemma 5.1, for any $\delta>0$ there exists $\hat{K}$ large enough such that
@@
$\sum_{i=1}^{t}(1-\hat{\rho})^{t-i}\hat{\xi}_{i}^{(k)}\leq\frac{\hat{K}}{\sqrt{n}}\sum_{i=1}^{t}(1-\hat{\rho})^{t-i}\big(1-{\hat{\rho}}/{2}\big)^{t-i}\big(1+\hat{\rho}\big)^{t-i}\leq\frac{C}{2\sqrt{n}},$
@@
$\Pr\Big(\sum_{i=1}^{t}\hat{\xi}_{i}^{(k)}(1-\hat{\rho})^{t-i}\geq\frac{C}{2\sqrt{n}}\Big)\leq\sum_{i=1}^{t}\Pr\Big(\hat{\xi}_{i}^{(k)}>\frac{\hat{r}_{i}}{\sqrt{n}}\Big)\leq\frac{\delta}{q}.$
Then, when $t\geq\frac{\log n}{2\log\big(\frac{q}{q-1}\big)}+\frac{\log(2/C)}{\log\big(\frac{q}{q-1}\big)}$, we have that $d_{t}^{(k)}\leq C/\sqrt{n}$ with probability at least $1-\delta/q$, and the result follows by a union bound over $k\in[q]$. ∎
@@ 6 Projection chain
Fix integers $q\geq 2$, $m\geq 0$ and a constant $L>0$. Let $X$ and $Y$ be two multinomial random variables with $m$ trials, where each trial leads to success of one of $q$ values with the probability of success for each value being $1/q$; i.e., $X,Y\sim\mathrm{Mult}(m;1/q,\ldots,1/q)$. Let $w=(w_{1},\dots,w_{q})\in\mathbb{Z}^{q}$ such that $\sum_{i=1}^{q}w_{i}=0$ and $\|w\|_{\infty}\leq L\sqrt{m}$. Then there exists a coupling of $X$ and $Y$ and a constant $c=c(q,L)>0$ independent of $m$ such that
$\Pr\big(X=Y+w\big)\ \geq\ c.$
@@
We couple $A_{t+1}$ and $A_{t+1}^{\prime}$ as follows. The percolation step is done independently in both copies. In the coloring step, first, the same random color $s$ is assigned to the largest connected component in both chains, and all other connected components with two or more vertices from both chains are assigned spins uniformly at random independently. In addition, in each $V_{i}$, if one copy has more isolated vertices than the other, the excess of isolated vertices in that copy are assigned spins independently. The coloring of the remaining isolated vertices will be coupled to fix the potential discrepancies created in each $V_{i}$ by this initial procedure as we described next.
For $i,j\in[q]$, let $\hat{a}_{ij}$, $\hat{a}_{ij}^{\prime}$ be the number of vertices assigned color $j$ in $V_{i}$ so far. We claim that there exists a constant $C>0$ such that with probability $\Omega(1)$ it holds for all $i,j\in[q]$ that
@@
To prove this, let $\{C_{k}\}_{k\geq 1}$ be the connected components in the first chain after the percolation step; assume $C_{1}$ is the largest component. Let $C(v)$ be the connected component of vertex $v$. Note that
$\sum_{k\geq 2}|C_{k}\cap V_{i}|^{2}=\sum_{k\geq 2}\sum_{v\in C_{k}\cap V_{i}}|C_{k}\cap V_{i}|=\sum_{k\geq 2}\sum_{v\in C_{k}\cap V_{i}}|C(v)\cap V_{i}|\leq\sum_{v\in V_{i}\setminus C_{1}}|C(v)|.$
@@
Let $\hat{n}_{i}$ be the number of vertices in $V_{i}$ excluding those in $C_{1}$ and the unassigned isolated vertices. If $C_{1}$ is assigned spin $s$, then conditioning on $\sum_{k\geq 2}|C_{k}\cap V_{i}|^{2}<100qA_{1}|V_{i}|$, by Hoeffding’s inequality, we deduce that for a large enough constant $A_{2}>0$, with probability at least $\frac{99}{100q^{2}}$, for any $j\in[q]$ with $j\neq s$, it holds that
@@
Similarly, we can get an analogous bound for $\hat{a}_{ij}^{\prime}$, and by a union bound (22) holds for all $i,j\in[q]$ such that $j\neq s$ with probability at least $96/100$.
By Lemma 2.3 and assumption A2, for a suitable constant $A_{3}>0$, we have for any $i\in[q]$ that
@@
Conditioning on this event and on the event that $\sum_{k\geq 2}|C_{k}\cap V_{i}|^{2}<100qA_{1}|V_{i}|$ both for all $i\in[q]$, we obtain via Hoeffding’s inequality that for a large enough constant $A_{2}^{\prime}>0$ that for any $i\in[q]$
$\Big|\hat{a}_{is}-|C_{1}\cap V_{i}|-\frac{\hat{n}_{s}}{q}\Big|$ $\leq(A_{2}^{\prime}+A_{3})\sqrt{n},~\text{and}$
$\Big|\hat{a}_{is}^{\prime}-|C_{1}^{\prime}\cap V_{i}|-\frac{\hat{n}_{s}}{q}\Big|$ $\leq(A_{2}^{\prime}+A_{3})\sqrt{n}$
with probability at least $99/(100q)$, which implies that (22) holds for all $i\in[q]$ and $j=s$ with that probability $99/100$ by a union bound.
@@
To formalize this, let $\mathcal{P}(r)$ be the set of pairs of configurations of the projection chain $A$ and $A^{\prime}$ such that any full configurations $X$ and $X^{\prime}$ that project onto $A$ or $A^{\prime}$, respectively, satisfy: $\|\alpha(X)-m\|_{\infty}\leq r/\sqrt{n}$, $\|\alpha(Y)-m\|_{\infty}\leq r/\sqrt{n}$, $\max_{k\in[q]}|\alpha_{1,k}(X)-\alpha_{1,k}(Y)|\leq r/\sqrt{n}$.
@@
For any $\delta\in(0,1)$, there exists $r=r(\delta)>0$ such that $\Pr(Z_{t+1}\in\mathcal{P}(r)\mid Z_{t}\in\mathcal{P}(r))\geq\delta$.
@@
Let $X_{t}$ and $Y_{t}$ be two configurations that project onto $Z_{t}=Z\in\mathcal{P}(r)$. Then, $\|\alpha(X_{t})-m\|_{\infty}\leq r/\sqrt{n}$. Let us assume that $\alpha_{1}(X_{t})\geq a$ (the case when $\alpha_{1}(X_{t})<a$ is analogous). By Fact 2.6, $F(\alpha_{1}(X_{t}))\leq F(a+r/\sqrt{n})$. Moreover, by Fact 2.5 and the mean value theorem, there exists a constant $\eta\in(0,1)$ such that $F(a+r/\sqrt{n})-a=\eta r/\sqrt{n}$ which implies that $(a+r/\sqrt{n})-F(a+r/\sqrt{n})=(1-\eta)r/\sqrt{n}$. In order for $\|\alpha(X_{t+1})-m\|_{\infty}\geq r/\sqrt{n}$, we would need $\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq\frac{(1-\eta)r}{\sqrt{n}}$. From Lemma 4.3, for any large enough constant $r$, we have
$\Pr\Big(\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq\frac{(1-\eta)r}{\sqrt{n}}\Big)\leq O(r^{-2}).$